OpenStudy (anonymous):

Use the concept of slope to find T such that the three points are collinear. (-6,1) (1,t)(10,5) Please explain! Totally lost

OpenStudy (anonymous):

\[ \frac{5-1}{10+6}=\frac{t-1}{1+6} \] Solve for t.

OpenStudy (anonymous):

\[ \frac{5-1}{10+6}=\frac{t-1}{1+6}\\ \frac{4}{16}=\frac 1 4=\frac{t-1}{1+6}\\ 7= 4 t -4\\ 4t =11\\ t=\frac{11}4 \]

OpenStudy (anonymous):

Thank you !!!

OpenStudy (anonymous):

What if it's switched? Like (t, -1)

OpenStudy (anonymous):

You switch the role ot t and -1

OpenStudy (anonymous):

of

OpenStudy (anonymous):

So you put me where?

OpenStudy (anonymous):

Them*

OpenStudy (anonymous):

Slope is \[ \frac{y_1-y_2}{x_1-x_2} \]

OpenStudy (anonymous):

(3,-3) (,-1)(8,6)

OpenStudy (anonymous):

T,-1

OpenStudy (anonymous):

You mimic my solution. You should be able to do that.

OpenStudy (anonymous):

There.